(Assume that 9 is the symbol for a partial derivative) For the equation (1/x)+(1/y)+(1/z)=1, evaluate 9z/9x and 9z/9y at the point (2,3,6)
1
Expert's answer
2020-10-20T16:58:21-0400

given expression is,

"\frac{1}{x}+\frac{1}{y}+\frac{1}{z}=1"


differentiate with respect to x

"\frac{-1}{x^2}-\frac{1}{z^2}\frac{\delta z}{\delta x} =0"



"\frac{\delta z}{\delta x}=-\frac{z^2}{x^2}"


"\frac{\delta z}{\delta x}_{(2,3,6)}=-\frac{6^2}{2^2}"

="-\frac{36}{4}=-9"


Now partially differentiate given expression with respect to y:

"\to" "\frac{-1}{y^2}-\frac{1}{z^2}\frac{\delta z}{\delta y} =0"


"\to \frac{\delta z}{\delta x} =\frac{-z^2}{y^2}"


"\to \frac{\delta z}{\delta x}_{(2,3,6)}=-\frac{6^2}{3^2}"


="-\frac{36}{9}"

=-4